OpenStudy (anonymous):

"f(x) = 5x + 2, g(x) = 2x - 1 Find (f o g)(x)" What does this mean?

OpenStudy (gorv):

it means f(g(x))

OpenStudy (gorv):

means value of g(x) is placed in function f

OpenStudy (gorv):

f(g(x))=f(2x-1)=5*(2x-1)+2

OpenStudy (gorv):

got it ??

Parth (parthkohli):

g(x) is NOT 2x - 1. It is given that 5x + 2g(x) = 2x - 1. Therefore,\[2g(x) = 2x - 1 - 5x = -3x - 1\]\[\iff g(x) = \dfrac{-3x-1}{2}\]

OpenStudy (anonymous):

Comment.

Parth (parthkohli):

French?

OpenStudy (gorv):

g=2x-1

Parth (parthkohli):

No, f(x) = 2x - 1.

OpenStudy (gorv):

i think you didnt read the Q properly

Parth (parthkohli):

@ganeshie8 Please confirm.

OpenStudy (anonymous):

It is a composite function so you plug in. SO: "f(x) = 5x + 2g(x) = 2x - 1 Find (f o g)(x)" What does this mean? It is saying that okay, you have two equations First is F(x) = 5x +2 And the second equation" G(x) = 2x-1 So that f circle g is F Of G of X so find the g equation substituted x for x. G(x)= 2x-1 you cant change because its x. then plug this is because it is the answer to plugging in x for x. SO F(G which is 2x-1) = 5(2x-1) +2 Do algebra. 10x-5+2 10x-3 is your answer:)

Parth (parthkohli):

Oh, wait.

Parth (parthkohli):

Looks like the question poster forgot a space or a line break.

ganeshie8 (ganeshie8):

lol yeah f(x) = 5x + 2g(x) = 2x - 1 technically gives f(x) = 2x-1 g(x) = (-3x-1)/2

Parth (parthkohli):

f(x) = 5x + 2, g(x) = 2x - 1

Parth (parthkohli):

^ that's probably what they meant

ganeshie8 (ganeshie8):

makes more sense, let me edit the question :)

OpenStudy (anonymous):

lololol